OpenStudy (anonymous):

Which term best describes the solution of the situation represented by the system of inequalities? (Assume that x >/= 0 and y >/= 0.) 3x + 4y

OpenStudy (zarkon):

do you mean \[x\ge 0\] \[y\ge 0\] \[3x+4y\le 8\] \[2x+4y\le 6\] \[f(x,y)=x+y\]